OpenStudy (anonymous):

A ladder leans on a wall and reaches 2.5m high. The foot of the ladder is 0.9m from the base of the wall. How long is the ladder? 5.

OpenStudy (anonymous):

2,5^2+0.9^2=x^2 6,25+0.81=x^2 x=sqrt(6,25+0.81) x=2,657... \[\approx\] 2,6

OpenStudy (anonymous):

Ignore the five

OpenStudy (anonymous):

Do you understand the pythagorean theorem?

OpenStudy (anonymous):

oh and, that should probably be rounded to 2,7*... not 2,6.